# BugSweeper: Function-Level Detection of Smart Contract Vulnerabilities Using Graph Neural Networks

### v1

@@ Abstract
Extended versionhttps://arxiv.org/abs/2406.12935
@@ Introduction
Blockchain is a decentralized and distributed ledger that enables open participation without third-party intermediaries and has attracted significant interest from academia and industry (krichen:blockchain). Ethereum extended blockchain capabilities by introducing smart contracts, which are digital agreements written in Solidity code that automatically execute transactions (buterin2013ethereum). However, these smart-contract programs introduce potential security vulnerabilities that can be exploited by malicious attackers. According to an empirical study (10.1145/3377811.3380364) that analyzed 47,368 smart contracts, many vulnerabilities, such as reentrancy and unchecked low-level calls, were reported. In particular, the DAO attack (TheDAO) exploited a reentrancy vulnerability to steal 3.6 million Ether (valued at \$60 million at the time). In addition, smart contracts cannot be modified once deployed, unlike traditional software. Fixing a deployed contract typically requires deleting the original and redeploying an updated version, which can be both inconvenient and costly. For these reasons, it is crucial to thoroughly verify the security of smart contracts before deployment.
@@
To overcome these drawbacks, researchers have increasingly leveraged deep learning models for smart contract vulnerability detection. For instance, Peculiar (wu2021peculiar) and ReVulDL (zhang2022reentrancy) utilize GraphCodeBERT (guographcodebert), while TMP (zhuang2021smart) and AME (liu2021smart) apply Graph Neural Networks (GNNs). These deep learning–based approaches can reduce analysis time and reliance on expert-crafted rules. Unfortunately, in a given smart contract, only a small fraction of the code is typically involved in a vulnerability. This observation motivates extracting vulnerability-specific code fragments for training the vulnerability detection model. Existing deep learning methods incorporate preprocessing steps for extraction. However, these preprocessing steps still depend on rigid, rule-based heuristics, resulting in several limitations:
@@
In this paper, we introduce BugSweeper, a novel graph-based framework for detecting vulnerabilities in smart contracts. BugSweeper employs a GNN trained on our proposed Function-Level Abstract Syntax Graph (FLAG) representation. It effectively identifies various vulnerability types in a fully automated and data-driven manner without relying on predefined expert rules, thereby overcoming the limitations of previous rule-based and preprocessing methods.
@@
Figure 1 provides an example of a reentrancy attack. In this scenario, the distributeExternal function uses call.value to send Ether to an external address, which in turn triggers the fallback function of a malicious contract. The fallback function—a default function that executes when no other function matches or when Ether is received—then repeatedly calls distributeExternal, causing it to re-enter and execute again before the previous invocation completes. Such cases illustrate that many security vulnerabilities in smart contracts arise from unsafe inter-function interactions. Inspired by this observation and motivated by the limitations of current rule-based preprocessing methods, we propose analyzing smart contracts at the function level for more precise vulnerability detection.
@@
  • A Novel Data-Driven, Multi-Class Vulnerability Detection Framework: We introduce BugSweeper, the first unified, rule-free framework that detects multiple vulnerability classes directly from code structure. Experiments conducted on three distinct vulnerability types—(i) Reentrancy, (ii) Unchecked Low-Level Calls, and (iii) Time Manipulation—demonstrate the effectiveness of our function-level, data-driven approach in generalizing to diverse vulnerabilities.
@@
Traditional methods, which rely on static analysis and symbolic execution, have been widely used to identify vulnerabilities in smart contracts. Specifically, static analysis–based methods detect vulnerability patterns by analyzing code, data flow, and control flow using predefined rules, without executing the program. Static analysis-based tools for smart contract vulnerability detection include SmartCheck (tikhomirov2018smartcheck) and Slither (feist2019slither). Symbolic execution-based methods identify vulnerabilities by tracking program execution paths and collecting symbolic values for conditions along those paths. Representative tools include Oyente (luu2016making), Osiris (torres2018osiris), Manticore (mossberg2019manticore), and Mythril (mueller2017mythril). Recent methods, such as Slise (wang2024efficiently), employ a hybrid approach that combines static analysis with symbolic execution.
@@
Various deep learning–based methods have been developed for vulnerability detection. TMP (zhuang2021smart) converts contracts into symbolic graphs, refines them by removing less important nodes, and normalizes the graph to apply Graph Neural Networks (GNNs). Furthermore, AME (liu2021smart) combines expert patterns with TMP. In contrast, Peculiar (wu2021peculiar) and ReVulDL (zhang2022reentrancy) focus on a contract’s data flow, targeting critical variables within a data-flow graph to analyze reentrancy vulnerabilities using the GraphCodeBERT model (guographcodebert). However, current deep learning approaches have not fully escaped the rigidity of earlier methods. Many still depend on rule-based components, such as the explicit expert patterns in AME or the heuristic-based graph simplification in TMP. This reliance on heuristics and specialization fundamentally limits their applicability to various vulnerabilities.
@@
In this section, we describe BugSweeper, a GNN-based framework for detecting smart contract vulnerabilities. We describe the model architecture in three parts: the Graph Constructor, Code Graph Neural Network, and Second-Stage Graph Neural Network. The overall architecture of the model is depicted in Figure 2.
@@
Previous research shows that AST-based approaches perform well on many code-related tasks (zhang2019novel), and FA-AST (wang2020detecting) proves that adding data-flow and control-flow details to an AST helps the model outperform a basic AST approach. Building on this, we use an abstract syntax graph that integrates data and control flow.
@@
Figure 3 illustrates how we construct FLAGs. First, we use the Solidity compiler solc as a parser to convert the source code into an AST. To enable function-level extraction, we mark the nodes in any subtree where the root is FunctionDefinition or VariableDeclaration. A basic AST fails to capture critical program execution details, so we augment it with supplementary edges to form a flow-augmented abstract syntax graph (FA-ASG). Specifically, we introduce control-flow and data-flow edges to create a richer semantic representation. Our FA-ASG consists of three distinct edge categories: the original structural edges from the AST, along with the newly added control-flow and data-flow edges, as detailed below.
@@
  • Control-flow edges show how the execution sequence flows in IfStatement, LoopStatement, and NextStatement nodes. For an IfStatement, CondTrue indicates the path when the condition is true, while CondFalse shows the path when it is false. A LoopStatement includes WhileExecution and ForExecution edges that point to the condition, and WhileNext and ForNext edges leading to the statements executed once the condition is met. Finally, NextStatement edges represent the sequential ordering of statements.
@@
For instance, a coverage of 1 includes only the subgraph of the target function itself. A coverage of 2 expands the graph to include the subgraphs of all directly called functions and variables (the 1-hop neighborhood). Extending this, a coverage of 3 further incorporates the subgraphs of functions and variables called by that second level (the 2-hop neighborhood). Through this iterative expansion, we generate enriched FLAGs that serve as our final units for analysis.
Figure 4: Solidity code is parsed into an AST and decomposed into function- and variable-level subgraphs. Increased coverage incorporates more inter-function connections. The graph for coverage1 stays the same across coverage levels, as it does not involve any function calls or variable references. Colors in the figure are solely for distinguishing the subgraphs.
@@
where $k$ is the layer number for message-passing depth $K$, $h^{0}_{v}$ is the initial node feature encoded via a BPE tokenizer, $\mathcal{N}(v)$ refers to the neighborhood function, which maps each node $v$ to its set of neighboring nodes in the graph. AGG aggregates feature vectors of neighboring nodes, and $\oplus$ represents the concatenation operation that merges the aggregated message with $h_{v}^{k-1}$. Table 1 shows the overall architecture of CGNN.
@@
We evaluate our proposed method, BugSweeper, against four baseline vulnerability detection tools—Slither (feist2019slither), SmartCheck (tikhomirov2018smartcheck), Mythril (mueller2017mythril), and Slise (wang2024efficiently)—as well as four deep learning–based techniques, TMP (zhuang2021smart), AME (liu2021smart), Peculiar (wu2021peculiar), and ReVulDL (zhang2022reentrancy).
Table 3 presents the results of our experiments, highlighting several important findings. First, traditional methods (Slither, SmartCheck, Mythril, Slise) exhibited poor performance in detecting reentrancy vulnerabilities, as evidenced by their low recall scores. Second, deep learning–based methods achieved significantly better performance overall. Notably, our function-level approach, BugSweeper, outperformed all other models on the AME dataset. BugSweeper demonstrated an improvement in precision (up to 99.87%) and achieved the highest overall F1-score (98.57%), outperforming the closest competitor by approximately 3.1%. These results indicate that BugSweeper significantly reduces both false positives and false negatives compared to existing approaches.
@@
BugSweeper w/o 2nd stage GNN (SAGE) TopKPool 87.74 72.48 78.71
@@
Table 5: Multi‑class vulnerability detection performance for different pooling methods on BugSweeper variants. Bolded results are the best and statistically significant.
@@ Acknowledgments
BugSweeper w/o 2nd stage GNN (SAGE) TopKPool 87.74 ($\pm$ 1.70) 72.48 ($\pm$ 1.00) 78.71 ($\pm$ 0.49) 0.786 0.013* 0.050*

### v2

@@ Abstract
Extended versionhttps://arxiv.org/abs/2512.09385
@@ Introduction
Blockchain is a decentralized and distributed ledger that enables open participation without third-party intermediaries and has attracted significant interest from academia and industry (krichen:blockchain). Ethereum extended blockchain capabilities by introducing smart contracts, which are digital agreements written in Solidity code that automatically execute transactions (buterin2013ethereum). However, these smart-contract programs introduce potential security vulnerabilities that can be exploited by malicious attackers. According to an empirical study (10.1145/3377811.3380364) that analyzed 47,368 smart contracts, many vulnerabilities, such as reentrancy and unchecked low-level calls, were reported. In particular, the DAO attack (TheDAO) exploited a reentrancy vulnerability to steal 3.6 million Ether (valued at \$60 million at the time). Furthermore, smart contracts cannot be modified once deployed, unlike traditional software. Fixing a deployed contract typically requires deleting the original and redeploying an updated version, which can be both inconvenient and costly. For these reasons, it is crucial to thoroughly verify the security of smart contracts before deployment.
@@
To overcome these drawbacks, researchers have increasingly leveraged deep learning models for smart contract vulnerability detection. For instance, Peculiar (wu2021peculiar) and ReVulDL (zhang2022reentrancy) utilize GraphCodeBERT (guographcodebert), while TMP (zhuang2021smart) and AME (liu2021smart) apply Graph Neural Networks (GNNs). These deep learning-based approaches can reduce analysis time and minimize reliance on expert-crafted rules. However, in a given smart contract, only a small fraction of the code is typically involved in a vulnerability. This observation motivates extracting vulnerability-specific code fragments for training vulnerability detection models. Existing deep learning methods incorporate preprocessing steps for extraction. However, these still depend on rigid, rule-based heuristics, resulting in several limitations:
@@
In this paper, we introduce BugSweeper, a novel graph-based framework for detecting vulnerabilities in smart contracts. BugSweeper employs a GNN trained on our proposed Function-Level Abstract Syntax Graph (FLAG) representation. It effectively identifies various vulnerability types in a fully automated and data-driven manner without relying on predefined expert rules, thereby overcoming the limitations of previous rule-based preprocessing methods.
@@
Figure 1 provides an example of a reentrancy attack. In this scenario, the distributeExternal function uses call.value() to send Ether to an external address, which in turn triggers the fallback function of a malicious contract. The fallback function—a default function that executes when no other function matches, or when Ether is received—then repeatedly calls distributeExternal, causing it to re-enter and execute again before the previous invocation completes. Such cases illustrate that many security vulnerabilities in smart contracts arise from unsafe inter-function interactions. Inspired by this observation and motivated by the limitations of current rule-based preprocessing methods, we propose analyzing smart contracts at the function level for more precise vulnerability detection.
@@
  • A Novel Data-Driven, Multi-Class Vulnerability Detection Framework: We introduce BugSweeper, the first unified and rule-free framework that detects multiple vulnerability classes directly from code structure. Experiments conducted on three distinct vulnerability types—(i) Reentrancy, (ii) Unchecked Low-Level Calls, and (iii) Time Manipulation—demonstrate the effectiveness of our function-level, data-driven approach that generalizes to diverse vulnerabilities.
@@
Traditional methods, which rely on static analysis and symbolic execution, have been widely used to identify vulnerabilities in smart contracts. Specifically, static analysis-based methods detect vulnerability patterns by analyzing code, data flow, and control flow using predefined rules, without executing the program. Static analysis-based tools for smart contract vulnerability detection include SmartCheck (tikhomirov2018smartcheck) and Slither (feist2019slither). Symbolic execution-based methods identify vulnerabilities by tracking program execution paths and collecting symbolic values for conditions along those paths. Representative tools include Oyente (luu2016making), Osiris (torres2018osiris), Manticore (mossberg2019manticore), and Mythril (mueller2017mythril). Recent methods, such as Slise (wang2024efficiently), employ a hybrid approach that combines static analysis with symbolic execution.
@@
Various deep learning-based methods have been developed for vulnerability detection. TMP (zhuang2021smart) converts contracts into symbolic graphs, refines them by removing less important nodes, and normalizes the graph to apply Graph Neural Networks (GNNs). Furthermore, AME (liu2021smart) combines expert patterns with TMP. In contrast, Peculiar (wu2021peculiar) and ReVulDL (zhang2022reentrancy) focus on a contract’s data flow, targeting critical variables within a data-flow graph to analyze reentrancy vulnerabilities using the GraphCodeBERT model (guographcodebert). However, current deep learning approaches have not fully escaped the rigidity of earlier methods. Many still depend on rule-based components, such as the explicit expert patterns in AME or the heuristic-based graph simplification in TMP. This reliance on heuristics and specialization fundamentally limits their applicability to various vulnerabilities.
@@
In this section, we introduce BugSweeper, a GNN-based framework for detecting smart contract vulnerabilities, and describe its architecture in three parts: the Graph Constructor, the Code Graph Neural Network, and the Second-Stage Graph Neural Network.
@@
Previous research showed that AST-based approaches perform well on many code-related tasks (zhang2019novel), and FA-AST (wang2020detecting) demonstrated that adding data-flow and control-flow details to an AST helps the model outperform a basic AST approach. Building on this, we use an abstract syntax graph that integrates data and control flow.
@@
Figure 3 illustrates how we construct FLAGs. First, we use the Solidity compiler solc as a parser to convert the source code into an AST. To enable function-level extraction, we mark the nodes in subtrees whose root is a FunctionDefinition or VariableDeclaration. A basic AST fails to capture critical program execution details, so we augment it with supplementary edges to form a flow-augmented abstract syntax graph (FA-ASG). Specifically, we introduce control-flow and data-flow edges to create a richer semantic representation. Our FA-ASG consists of three distinct edge categories: the original structural edges from the AST, along with the newly added control-flow and data-flow edges, as detailed below.
@@
  • Control-flow edges show how the execution sequence flows in IfStatement, LoopStatement, and NextStatement nodes. For an IfStatement, CondTrue indicates the path when the condition is true, while CondFalse shows the path when it is false. A LoopStatement includes WhileExecution and ForExecution edges that point to the condition, and WhileNext and ForNext edges leading to the statements executed once the condition is met. NextStatement edges represent the sequential ordering of statements.
@@
For instance, a coverage of 1 includes only the subgraph of the target function itself. A coverage of 2 expands the graph to include the subgraphs of all directly called functions and variables (the 1-hop neighborhood). At a coverage of 3, the graph further incorporates the subgraphs of functions and variables called by that second level (the 2-hop neighborhood). Through this iterative expansion, we generate enriched FLAGs that serve as our final units for analysis.
Figure 4: Solidity code is parsed into an AST and decomposed into function- and variable-level subgraphs. Higher coverage incorporates more inter-function connections. The graph for coverage1 stays the same across coverage levels, as it does not involve any function calls or variable references. Colors in the figure are solely for distinguishing the subgraphs.
@@
where $k$ is the layer number for message-passing depth $K$, $h^{0}_{v}$ is the initial node feature encoded via a BPE tokenizer, and $\mathcal{N}(v)$ refers to the neighborhood function, which maps each node $v$ to its set of neighboring nodes in the graph. AGG aggregates feature vectors of neighboring nodes, and $\oplus$ represents the concatenation operation that merges the aggregated message with $h_{v}^{k-1}$. Table 1 shows the overall architecture of CGNN.
@@
We evaluate our proposed method, BugSweeper, against four baseline vulnerability detection tools—Slither (feist2019slither), SmartCheck (tikhomirov2018smartcheck), Mythril (mueller2017mythril), and Slise (wang2024efficiently)—as well as four deep learning-based techniques, TMP (zhuang2021smart), AME (liu2021smart), Peculiar (wu2021peculiar), and ReVulDL (zhang2022reentrancy).
Table 3 presents the results of our experiments, highlighting several important findings. First, traditional methods (Slither, SmartCheck, Mythril, Slise) exhibited poor performance in detecting reentrancy vulnerabilities, as evidenced by their low recall scores. Second, deep learning-based methods achieved significantly better performance overall. Notably, our function-level approach, BugSweeper, outperformed all other models on the AME dataset. BugSweeper demonstrated an improvement in precision (up to 99.87%) and achieved the highest overall F1-score (98.57%), outperforming the closest competitor by approximately 3.1%. These results indicate that BugSweeper significantly reduces both false positives and false negatives compared to existing approaches.
@@
BugSweeper (CGNN only) TopKPool 87.74 72.48 78.71
@@
Table 5: Multi‑class vulnerability detection performance for different pooling methods on BugSweeper variants. BugSweeper (CGNN only) denotes a single-stage CGNN with GraphSAGE. Bolded results are the best and statistically significant.
@@ Acknowledgments
BugSweeper w/o second-stage GNN (SAGE) TopKPool 87.74 ($\pm$ 1.70) 72.48 ($\pm$ 1.00) 78.71 ($\pm$ 0.49) 0.786 0.013* 0.050*